OpenStudy (jackman):

Two car washers, Lydia and Jackie, are working on your car. Lydia can complete the work in 2 hours, while Jackie can complete the work in 1 hour. How many hours do the car washers take if they work together?

OpenStudy (ashlyn_loves_cats):

an hour and 30 mins

OpenStudy (jackman):

these are the options 1/2 3/2 2/3 4/3

OpenStudy (mathstudent55):

If Jackie can complete the job alone in 1 hour, then when both work together, they must take less than 1 hour.

OpenStudy (mathstudent55):

Jackie completes the job in 1 hour. In 1 hour, Jackie does 1/1 = 1 full job Lydia completes the job in 2 hours. In 1 hour, Lydia completes 1/2 of the job. Working together, they take x time to complete the job. In 1 hour, they do 1/x of the job. \(\dfrac{1}{1} + \dfrac{1}{2} = \dfrac{1}{x} \) Solve for x.